OpenStudy (anonymous):

whats 5,602-2,344 times 4

OpenStudy (lizz123):

The answer is: 13,032!! By the way: use can use the calculator

OpenStudy (anonymous):

5602-2344=3258

OpenStudy (lizz123):

3258*4=13,032

OpenStudy (anonymous):

umm... lizz123, your not supposed to give answers.

OpenStudy (anonymous):

Just help...

OpenStudy (lizz123):

@Aaleelee123456

OpenStudy (anonymous):

First, you want to do the multiplications by 4. Next, subtract that answer from 5602. Think order of operations.

OpenStudy (anonymous):

its easy use a calculator in your laptop or use a google calculator

OpenStudy (anonymous):

Easy! Just take your problem 5,602 simplify by 2,344 Your answer should be 3,258 Now take your answer : 3,258 And multiply by 4 3,258 x 4 = 13,032 So your answer is 13,032! Hope this helped (:

OpenStudy (anonymous):

hereis the link to the google calulator

sammixboo (sammixboo):

Yes, follow the Order of Operations! P - Do the Parenthesis first E - Then the exponents M/D - now you are supposed to do multiplication or division starting from the left of the equations and ending at the right A/S - Lastly, you do addition or subtraction starting frpm the left of the equations and ending at the right

OpenStudy (anonymous):

Simply, the answer is -3774, if the equation is \[5602 - 2344 \times 4\] it cannot be \[5602 - 2344(4)\] or else the answer will be different.